OpenStudy (anonymous):

please help me verify tan(2 pi-x)=-tanx using the difference formula for tan

OpenStudy (jdoe0001):

\(\bf tan(2\pi -x) = -tan(x)\\ \quad \\ \quad \\ tan(2\pi -x) =\cfrac{tan(2\pi)-tan(x)}{1+tan(2\pi)tan(x)}\\ \quad \\ \textit{so, what's}\quad tan(2\pi)\quad ? \)

OpenStudy (anonymous):

Wow ok! I must have just been over thinking it. Thanks!